Find three consecutive positive even integers such that the product of the second and third
integers is twenty more than ten times the first integer.

Answer by nyc_function(2741)

Let x = first integer
Let x + 2 = second integer
Let x + 4 = third integer
I will give you the equation and you can do the work.
(x + 2)(x + 4) = 10x + 20